# A method to derive material-specific spin-bath model descriptions of materials displaying prevalent spin physics (for simulation on NISQ devices)

### v2

@@ Abstract
Magnetism and spin physics are true quantum mechanical effects and their description usually requires multi reference methods and is often hidden in the standard description of molecules in quantum chemistry. In this work we present a twofold approach to the description of spin physics in molecules and solids. First, we present a method that identifies the single-particle basis in which a given subset of the orbitals is equivalent to spin degrees of freedom for models and materials which feature significant spin physics at low energies. We introduce a metric for the spin-like character of a basis orbital, of which the optimization yields the basis containing the optimum spin-like basis orbitals. Second, we demonstrate an extended Schrieffer-Wolff transformation method to derive the effective Hamiltonian acting on the subspace of the Hilbert space in which the charge degree of freedom of electron densities in the spin-like orbitals is integrated out. The method then yields an effective spin-bath Hamiltonian description for the system. This extended Schrieffer-Wolff transformation is applicable to a wide range of Hamiltonians and has been utilized in this work for model Hamiltonians as well as the active space Hamiltonian of molecular chromium bromide.
@@ I Introduction
The theoretical study of the physics of many-body quantum system has remained notoriously difficult on conventional computers not least due to the intractable exponential growth of the corresponding Hilbert spaces. This holds true already for simple model Hamiltonians, but becomes even more problematic for the study of real materials due to the vast number of degrees of freedom and interactions. To investigate certain aspects of these materials it is therefore imperative to either utilize significant approximations or to obtain a much more concise Hamiltonian focused on accurately describing this aspect. In this article, we concentrate on the description of the spin physics of actual materials. The spin physics is relevant for the understanding of the material’s magnetic properties [1] as well as their experimental identification with spin resonance spectroscopy techniques [2]. The quantum simulation of spin system Hamiltonians is also perfectly suited as an application of quantum computing in general [3] and analogue quantum computing in particular [4, 5, 6].
@@
We present a method for the derivation of the effective Hamiltonian description for the relevant spin degrees of freedom of an actual material embedded in an environment of fermions. The first part of the method entails the identification of potential relevant orbital spin degrees of freedom. We propose a new metric for the classification of orbital spin degrees of freedom and present a systematic way to determine the orbital basis that contains the best realizations of these spins, see also  [7]. Given the presence of relevant spins, we proceed with the second part of the method, a Schrieffer-Wolff transformation procedure [8, 9] aimed at decoupling the subspace of the Hilbert space containing these spin degrees of freedom from the remaining high-energy subspaces. Many different Schrieffer-Wolff and related unitary transformation schemes have been proposed, comprising perturbative [8, 10, 11], variational [12], and continuous [13, 14, 15] variants. So far, the application of these methods has been limited to concise model Hamiltonians. We present an extended version of the perturbative Schrieffer-Wolff transformation that can be applied to generic Hamiltonian descriptions of materials, given the presence of relevant low-energy spin physics in the material. We demonstrate the accuracy of the effective spin-bath model Hamiltonians resulting from our Schrieffer-Wolff transformation for the description of the low-energy physics both for model Hamiltonians with established results [8, 16] as well as an actual material. We further discuss the extent to which the quality of the effective spin-bath model Hamiltonian derived with our proposed method can be predicted.
@@ II Methodology
In Section II.1 we define the notion of spin-like orbitals and in Section II.2 we propose the local parity as a metric for the spin-like character of orbitals. In Section II.3 we present our parity optimization procedure as a way to determine the spin-like orbitals of a system. We discuss the basic concept of the Schrieffer-Wolff transformation in Section II.4. We detail our proposed extended Schrieffer-Wolff transformation method in Sections II.5 and II.6. In Section II.7 we show how the described methods are combined into our workflow for deriving effective spin-bath model Hamiltonians for materials with relevant low-energy spin physics.
@@ II.1 Spin-like orbitals
We consider an orbital $\phi_{i}$ as being spin-like only if the electron density $n_{i}$ contained in it is strictly equal to one. This requires that the average electron density in the orbitals $\phi_{i}$ satisfies $\langle n_{i}\rangle=1$. It furthermore requires that the fluctuations around this average electron density satisfy $\delta n_{i}\rightarrow 0$. Negligible fluctuations around the average electron density imply that the electron density of the orbital does not affect the low-energy dynamics of the system and vice versa. Just an average electron density $\langle n_{i}\rangle=1$ places no restrictions on the orientation and dynamics of the electron spin in the orbital $\phi_{i}$. When both requirements are simultaneously met by the states in the low-energy Hilbert space, the dynamics of the electron density in the orbital $\phi_{i}$, often referred to as the charge degree of freedom, becomes superfluous to the description of the dynamics of the system. The electron density in the orbital $\phi_{i}$ consequently couples to the remainder of the system exclusively via its spin degree of freedom. It is then sufficient to represent the degrees of freedom of the electron density contained in the orbital $\phi_{i}$ as a pure spin degree of freedom and to employ the associated spin operator algebra. The local Hilbert space $\mathcal{H}_{i}$ of a spin degree of freedom is half the size of the local Hilbert space of a fermionic orbital and is furthermore naturally represented on a Qubit. We therefore aim to determine each spin-like basis orbital or linear combinations thereof meeting both stated requirements, so that they can be represented as spins.
@@ II.2 Local parity
We propose the ground state local parity $P_{i}$ as measure for the spin-like character of an orbital $\phi_{i}$. The operator representation of the local parity reads
@@
[…]
@@
with $02\geq\alpha>$ 0, because the resulting fluctuations in the electron density $\delta n_{i}\neq 0$ manifest themselves in strictly positive contributions to the local parity. An alternative and more useful operator representation of the local parity reads
@@
denotes the two-electron reduced density matrix (2-RDM) respectively. If we choose $|n\rangle=|\psi_{0}\rangle$, with $|\psi_{0}\rangle$ a good approximation of the ground state of the system, we can identify orbitals $\phi_{i}$ for which $\langle P_{i}\rangle_{0}=\langle\psi_{0}|P_{i}|\psi_{0}\rangle=-1+\varepsilon$, with $\varepsilon\rightarrow 0$, as spin-like orbitals of the system. In general, the spin-like orbitals of the system do not coincide with the basis orbitals. We therefore require a method to determine the set $\{\phi_{i}\}$ of orthonormal linear combinations of basis orbitals, for which the local parities most closely approach $\{p_{i}\}=-1$.
@@ II.3 Local parity optimization of the orbital basis
with the same rotation being performed for the Hermitian conjugates of the operators. From the reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ we can compute the local parity of an orbital $\phi_{q}$, which results from the linear combination of orbitals $\phi_{i}$ and $\phi_{j}$, as $\langle P_{q}\rangle_{0}(\theta)$, where $\langle\dots\rangle_{0}$ denotes the expectation value with respect to the ground state. The local parity $\langle P_{q}\rangle_{0}(\theta)$ is an analytic, $2\pi$-periodic function of the rotation angle $\theta$. We find the extremal points $\theta_{n}$ of the function $\langle P_{q}\rangle_{0}(\theta)$ in the domain $\theta\in[0,2\pi)$ from
@@
The analytic expression for the derivatives of the function $\langle P_{q}\rangle_{0}(\theta)$ are shown in appendix A. If a solution $\theta_{m}$ exists which satisfies $d^{2}\langle P_{q}\rangle_{0}/d\theta^{2}|_{\theta_{m}}>0$ and $\langle P_{q}\rangle_{0}(\theta_{m})\leq\langle P_{q}\rangle_{0}(\theta_{l})\,\forall\theta_{l}\in\theta_{n}$, we accept the rotation attempt $i\rightarrow q$ and $j\rightarrow p$ with rotation angle $\theta_{m}$ and we reject the rotation attempt otherwise. By repeating the procedure for each pair of basis orbitals $\phi_{i}$ and $\phi_{j}$, we arrive at at an orthonormal basis in which the local parities have taken up extremal values. We then identify the orbitals $\phi_{q}$ of the resulting basis for which $\langle P_{q}\rangle_{0}=-1+\varepsilon$, with $\varepsilon$ an arbitrary small, positive value, as the spin-like orbitals of the system. Basis orbitals $\phi_{p}$ with a local parity $\langle P_{p}\rangle_{0}\simeq+1$ can be regarded as beneficial for the purpose of separating the system’s spin degrees of freedom from their respective environment since they experience exclusively a transfer of an even number of particles only, such that the spin degree of freedom of electrons occupying the orbitals $\phi_{p}$ becomes insignificant. In addition, we have implemented an optimization scheme using the full gradient. Currently, the pairwise optimization is typically more efficient.
@@ II.4 Perturbative similarity transformation
[…]
@@
where the terms originating from $[S,\,V]/2$ contain the perturbative corrections arising from the consecutive application of two block-offdiagonal operators. A subsequent projection to the subspaces
@@ II.5 Symmetry specification of block-offdiagonal operators
$=$ $\sum_{\left\{v\right\}}\alpha_{v}\left[\left(\prod_{i,j}o^{j}\sum_{q(i)}x^{i}_{q(i)}\right)+\left(\prod_{i,j}o^{j}\sum_{q(i)}x^{i}_{q(i)}\right)^{\dagger}\right]\,,$
@@
since $\mathcal{V}_{1}$ includes additional unique operator sequences generated by $[H_{0},\hat{e}_{v}]$. In the vector spaces $\mathcal{V}$ we can translate equation (20) for the generator $S$ of the Schrieffer-Wolff transformation to
@@ II.7 Full workflow
We start with an ab-initio electronic structure calculation [18] of the material to determine its ground state. The electronic structure method needs to be a post-Hartree-Fock or related method, this excludes density functional theory, to capture the effect of correlations in the two-particle reduced density matrix $\rho^{(2)}$. From the electronic structure calculation we obtain the single particle basis orbitals for the one-electron and two-electron integrals which specify the Hamiltonian description. For the ground state of the calculation we compute the one-particle and two-particle reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$.
@@
We utilize the reduced density matrices to assign a local parity $\langle P_{i}\rangle_{0}$ to the basis orbitals $\phi_{i}$. We then perform pairwise rotations of the basis orbitals to determine the basis in which the local parities of the basis orbitals is extremized. If there exist optimized basis orbitals $\phi_{q}$ with $\langle P_{q}\rangle_{0}+1<\varepsilon$, where we typically chose $\varepsilon\leq 10^{-1}$, we proceed with the subsequent steps of the workflow. If no spin-like orbitals are found, we terminate the workflow.
@@ III.1 Spin-bath chain model
Figure 1: Schematic representation of the lattice model described by the Hamiltonian  (53). The lattice sites $A_{i}$ are shown in grey. On the orange lattice sites $B_{i}$ the fermions experience a strong repulsive Hubbard interaction $U$ which energetically discourages $n_{B_{i}}\neq 1$. The hybridization between the lattice sites $A_{i}$ and $A_{i+1}$ is given by $t<0$. We impose periodic boundary conditions via a hybridization $t$ between the lattice sites $A_{N=4}$ and $A_{1}$. There is a small hybridization $0>V>t$ between lattice sites $A_{i}$ and $B_{i}$. The local Hubbard density-density interaction of strength $U\gg|t|$ is strongly repulsive. The connectivity and the coupling constants of the lattice models are chosen such that the lattice sites $B_{i}$ should be good realizations of spin-like orbitals.
@@
[…]
@@
where $a^{\dagger}_{i\sigma}$ creates a fermion of spin $\sigma$ on lattice site $A_{i}$ and $b^{\dagger}_{i\sigma}$ creates a fermion of spin $\sigma$ on lattice site $B_{i}$. We impose periodic boundary conditions for the lattice sites $A_{i}$. A graphic representation of the lattice model is displayed in Fig. 1. The coupling constants satisfy $U\gg|t|>|V|$ and $0>V>t$. The repulsive Hubbard interaction on lattice sites $B_{i}$ in combination with the weak hybridization between $B_{i}$ and $A_{i}$ places a significant energy penalty on $n_{B_{i}}\neq 1$. Low-energy eigenstates of the Hamiltonian should therefore satisfy $n_{B_{i}}\equiv 1$ and the lattice sites $B_{i}$ be considered good realizations of spin-like orbitals.
We have calculated the low energy spectrum of the Hamiltonian of a chain of $M=4$ unit cells containing $N=8$ fermions using a numerical diagonalization method. The reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ have been computed for the corresponding ground state $|\psi_{0}\rangle$. The eigenstates of $\rho^{(1)}$ form the natural orbital basis. We find eigenvalues $\langle n_{i}\rangle=1$ of $\rho^{(1)}$ for $N_{\phi_{i}}=2$ natural basis orbital $\phi_{i}$.
Figure 2: Local parity $\langle P_{q}\rangle_{0}$ of the basis orbitals of (a) the natural orbitals basis and (b) the parity optimized basis of the spin-bath chain model described by eq. (53). (a) The dark grey line shows the local parities of the original basis orbitals and the dashed orange line indicates the local parities of the natural basis orbitals. The blue circles display the average electron density $\langle n_{q}\rangle_{0}$ in the natural basis orbitals. In the natural basis we find $N_{\phi_{q}}=2$ orbitals $\phi_{q}$ for which $\langle n_{q}\rangle_{0}=1$, but the respective values of the local parities $\langle P_{q}\rangle_{0}>0$ highlight that $|\delta n_{q}|\gg 0$. (b) The dashed orange line now displays the local parities of the parity optimized basis orbitals. The blue circles show the average electron density $\langle n_{q}\rangle_{0}$ in these optimized basis orbitals. In the optimized basis we find $N_{\phi_{q}}\geq 4$ orbitals $\phi_{q}$ with local parity $\langle P_{q}\rangle_{0}\simeq-1$. We also observe $\langle n_{q}\rangle_{0}=1$ for each orbital. The orbitals $\phi_{q\leq 3}$ of the optimized basis are considered spin-like.

### v3

# A method to derive material-specific model descriptions of spin degrees of freedom coupled to a fermionic bath for materials displaying prevalent spin physics

@@ Abstract
Magnetism and spin physics are true quantum mechanical effects and their description usually requires multi reference methods and is often hidden in the standard description of molecules in quantum chemistry. In this work we present a twofold approach to the description of spin physics in molecules and solids. First, we present a method that identifies the single-particle basis in which a given subset of the orbitals is equivalent to spin degrees of freedom for models and materials which feature significant spin physics at low energies. We introduce a metric for the spin-like character of a basis orbital, of which the optimization yields the basis containing the optimum spin-like basis orbitals. Second, we demonstrate an extended Schrieffer-Wolff transformation method to derive the effective Hamiltonian acting on the subspace of the Hilbert space in which the charge degree of freedom of electron densities in the spin-like orbitals is integrated out. The method then yields an effective Hamiltonian describing spins coupled to a fermionic environment. This extended Schrieffer-Wolff transformation is applicable to a wide range of Hamiltonians and has been utilized in this work for model Hamiltonians as well as the Hamiltonian describing the active orbital space of molecular chromium bromide. This is achieved by reformulating the highly non-linear Schrieffer-Wolff equations into a linear set of equations corresponding to an operator basis.
@@ I Introduction
The theoretical study of the physics of many-body quantum systems has remained notoriously difficult on conventional computers not least due to the intractable exponential growth of the corresponding Hilbert spaces. This holds true already for simple model Hamiltonians, but becomes even more problematic for the study of real materials due to the vast number of degrees of freedom and interactions. To investigate certain aspects of these materials it is therefore imperative to either utilize significant approximations or to obtain a much more concise Hamiltonian focused on accurately describing this aspect. In this article, we concentrate on the description of the spin physics of actual materials. The spin physics is relevant for the understanding of the material’s magnetic properties [1] as well as their experimental identification with spin resonance spectroscopy techniques [2, 3]. The quantum simulation of spin system Hamiltonians is also perfectly suited as an application of quantum computing in general [4] and analogue quantum computing in particular[5].
@@
We present a method for the derivation of the effective Hamiltonian description for the relevant spin degrees of freedom of an actual material embedded in an environment of fermions. The first part of the method entails the identification of potential relevant orbital spin degrees of freedom. We propose a new metric for the classification of orbital spin degrees of freedom and present a systematic way to determine the orbital basis that contains the best realizations of these spins, see also  [6]. Given the presence of relevant spins, we proceed with the second part of the method, a Schrieffer-Wolff transformation procedure [7, 8] aimed at decoupling the subspace of the Hilbert space containing these spin degrees of freedom from the remaining high-energy subspaces containing the charge degrees of freedom. Many different Schrieffer-Wolff and related unitary transformation schemes have been proposed, comprising perturbative [7, 9, 10], variational [11], and continuous [12, 13, 14] variants. So far, the application of these methods has been limited to concise model Hamiltonians. We present an extended version of the perturbative single step Schrieffer-Wolff transformation that can be applied to generic Hamiltonian descriptions of materials, given the presence of relevant low-energy spin physics in the material. By restricting ourselves to spin-like orbitals that have by definition no or only small charge fluctuations we ensure that a perturbative treatment of the charge fluctuations is justified. This is also a main difference to canonical transformation approaches [15] where on tries achieve a canonical diagonalization. We restrict our canonical transformation to a perturbative regime where the application does not create large higher order terms.
We demonstrate the accuracy of the effective spin-bath model Hamiltonians, where the spin-like orbitals are treated as spins coupled to a fermionic bath, resulting from our Schrieffer-Wolff transformation for the description of the low-energy physics both for simple model Hamiltonians with established results [7, 16] as well as actual materials. We further discuss the extent to which the quality of the effective model Hamiltonian for the spins coupled to the fermionic environment derived with the proposed method can be anticipated.
We would like to point out that we do start with an approximate multi reference calculation to determine a suitable single particle basis. However, the Schrieffer-Wolff expansion, as explained below, starts from the full quantum chemical description only using the adapted basis set from the multi reference calculation. At first glance this may appear as a waste of resources. However, as shown below, already starting in a suitable single particle basis leads to a reasonably good spin-bath model, which then improved by the Schrieffer Wolff transformation. This approach is in strong contrast to a traditional setup, where one tries to extract either an effective Hubbard model using DFT+U [17, 18], DFT+DMFT [19, 20, 21] or constrained RPA [22, 23, 24, 25, 26] in the framework of maximally localized Wannier functions [27, 28, 29], which is can then be transformed into a spin-math model. Or alternatively using DFT based approaches using the magnetic force theorem [30, 31]. In these approaches various approximation are staggered. In our work we only rely on the Schrieffer-Wolff transformation.
@@ II Methodology
In Section II.1 we define the notion of spin-like orbitals and in Section II.2 we propose the local parity as a metric for the spin-like character of orbitals. In Section II.3 we present our parity optimization procedure as a way to determine spin-like orbitals, as well as a way to gauge the presence of relevant low energy spin physics in the system. We discuss the basic concept of the Schrieffer-Wolff transformation in Section II.4. We provide our proposed extended Schrieffer-Wolff transformation method in detail in Sections II.5 and II.6. In Section II.7 we show how the described methods are combined into our workflow for deriving effective model Hamiltonians of spins in a fermionic environment for materials with relevant low-energy spin physics.
@@ II.1 Spin-like orbitals
We consider an orbital $\phi_{i}$ as being spin-like only if the electron density $n_{i}$ contained in it is strictly equal to one. This requires that the average electron density in the orbitals $\phi_{i}$ satisfies $\langle n_{i}\rangle=1$. It furthermore requires that the fluctuations around this average electron density satisfy $\delta n_{i}\rightarrow 0$. Negligible fluctuations around the average electron density imply that electron density of the orbital does not affect the low-energy dynamics of the system and vice versa. Just an average electron density $\langle n_{i}\rangle=1$ places no restrictions on the orientation and dynamics of the electron spin in the orbital $\phi_{i}$. When both requirements are simultaneously met by the states in the low-energy Hilbert space, the dynamics of the electron density in the orbital $\phi_{i}$, often referred to as the charge degree of freedom, becomes superfluous to the description of the dynamics of the system. The electron density in the orbital $\phi_{i}$ consequently couples to the remainder of the system exclusively via its spin degree of freedom. It is then sufficient to represent the degrees of freedom of the electron density contained in the orbital $\phi_{i}$ as a pure spin degree of freedom and to employ the associated spin operator algebra. The local Hilbert space $\mathcal{H}_{i}$ of a spin degree of freedom is half the size of the local Hilbert space of a fermionic orbital and is furthermore naturally represented on a Qubit. We therefore aim to determine each spin-like basis orbital or linear combinations thereof meeting both stated requirements, so that they can be represented as spins.
@@ II.2 Local parity
We propose the ground state local parity $P_{i}$ as a measure for the spin-like character of an orbital $\phi_{i}$. The operator representation of the local parity reads
@@
$\left\{|0\rangle,\,|\uparrow\rangle,\,|\downarrow\rangle,\,|\uparrow\downarrow\rangle\r$ight\}\,, (2)
@@
with $2\geq\alpha>0$, because the resulting fluctuations in the electron density $\delta n_{i}\neq 0$ manifest themselves in strictly positive contributions to the local parity. An alternative and more useful operator representation of the local parity reads
@@
denotes the two-electron reduced density matrix (2-RDM) respectively. If we choose $|n\rangle=|\psi_{0}\rangle$, with $|\psi_{0}\rangle$ a good approximation of the ground state of the system, we can identify orbitals $\phi_{i}$ for which $\langle P_{i}\rangle_{0}=\langle\psi_{0}|P_{i}|\psi_{0}\rangle=-1+\varepsilon$, with $\varepsilon\rightarrow 0$, as spin-like orbitals of the system. In general, the spin-like orbitals of the system do not coincide with the basis orbitals typically used in standard basis sets of quantum chemistry codes. We therefore require a method to determine the set $\{\phi_{i}\}$ of orthonormal linear combinations of basis orbitals, for which the local parities most closely approach $\{p_{i}\}=-1$.
@@ II.3 Local parity optimization of the orbital basis
with the same rotation being performed for the Hermitian conjugates of the operators. From the reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ we can compute the local parity of an orbital $\phi_{q}$, which results from the linear combination of orbitals $\phi_{i}$ and $\phi_{j}$, as $\langle P_{q}\rangle_{0}(\theta)$. The local parity $\langle P_{q}\rangle_{0}(\theta)$ is an analytic, $2\pi$-periodic function of the rotation angle $\theta$. We find the extremal points $\theta_{n}$ of the function $\langle P_{q}\rangle_{0}(\theta)$ in the domain $\theta\in[0,2\pi)$ from
@@
The analytic expression for the derivatives of the function $\langle P_{q}\rangle_{0}(\theta)$ are shown in appendix A. If a solution $\theta_{m}$ exists which satisfies $d^{2}\langle P_{q}\rangle_{0}/d\theta^{2}|_{\theta_{m}}>0$ and $\langle P_{q}\rangle_{0}(\theta_{m})\leq\langle P_{q}\rangle_{0}(\theta_{l})\,\forall\theta_{l}\in\theta_{n}$, we accept the rotation attempt $i\rightarrow q$ and $j\rightarrow p$ with rotation angle $\theta_{m}$ and we reject the rotation attempt otherwise. By repeating the procedure for each pair of basis orbitals $\phi_{i}$ and $\phi_{j}$, we arrive at at an orthonormal basis in which the local parities have taken up extremal values. We then identify the orbitals $\phi_{q}$ of the resulting basis for which $\langle P_{q}\rangle_{0}=-1+\varepsilon$, with $\varepsilon$ an arbitrary small, positive value, as the spin-like orbitals of the system. Basis orbitals $\phi_{p}$ with a local parity $\langle P_{p}\rangle_{0}\simeq+1$ can be regarded as beneficial for the purpose of separating the system’s spin degrees of freedom from their respective environment since they experience exclusively a transfer of an even number of particles only, such that the spin degree of freedom of electrons occupying the orbitals $\phi_{p}$ becomes insignificant. If the optimization gets stuck we found that first optimizing for minimal parities, and then optimizing the non spin-like orbitals for maximal parity we get out off local minima. In addition, we have implemented an optimization scheme using the full gradient. Currently, the pairwise optimization is typically more efficient, but this approach may help in the future to improve convergence.
@@ II.4 Perturbative similarity transformation
where for generators $S$ satisfying $\|S\|\ll\|H_{0}\|$, with $\|\cdot\|$ a suitable norm, one can approximate the expression as
@@
where the terms originating from [S,V]/2[S,\,V]/2 contain the perturbative corrections arising from the consecutive application of two block-offdiagonal operators. A subsequent projection to the subspaces
@@ II.5 Symmetry specification of block-offdiagonal operators
[…]
@@
since $\mathcal{V}_{1}$ includes additional unique operator sequences generated by $H_{0},\hat{e}_{v}]$. In the vector spaces $\mathcal{V}$ we can translate equation (20) for the generator $S$ of the Schrieffer-Wolff transformation to
@@
In summary, we have mapped the non-linear Schrieffer–Wolff equations, as presented in the original work, [7] to a linear set of equations, see also Sec. III.3. Without this representation we could not apply the Schrieffer–Wolff approach to quantum chemistry descriptions including four-index terms.
@@ II.7 Full workflow
We start with an ab-initio electronic structure calculation [33] of the material to determine its ground state. The electronic structure method needs to be a post-Hartree-Fock or a related method, this excludes density functional theory, to capture the effect of correlations in the two-particle reduced density matrix $\rho^{(2)}$. From the electronic structure calculation we obtain the single particle basis orbitals and the corresponding one-electron and two-electron integrals which specify the Hamiltonian description. For the ground state of the calculation we compute the one-particle and two-particle reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$.
@@
We utilize the reduced density matrices to assign a local parity $\langle P_{i}\rangle_{0}$ to the basis orbitals $\phi_{i}$. We then perform pairwise rotations of the basis orbitals to determine the basis in which the local parities of the basis orbitals is extremized. If there exist optimized basis orbitals $\phi_{q}$ with $\langle P_{q}\rangle_{0}+1<\varepsilon$, where we typically chose $\varepsilon\leq 10^{-1}$, we proceed with the subsequent steps of the workflow. If no spin-like orbitals are found, we terminate the workflow, because the absence of spin-like orbitals also indicates the absence of relevant spin physics in the system under consideration.
@@ III.1 Spin-bath chain model
Figure 1: Schematic representation of the lattice model described by the Hamiltonian  (53). The lattice sites $A_{i}$ are shown in grey. On the orange lattice sites $B_{i}$ the fermions experience a strong repulsive Hubbard interaction $U$ which energetically discourages $n_{B_{i}}\neq 1$. We refer to the lattice sites $A_{i}$ and $B_{i}$ as the initial basis of the system. The hybridization between the lattice sites $A_{i}$ and $A_{i+1}$ is given by $t<0$. We impose periodic boundary conditions via a hybridization $t$ between the lattice sites $A_{N=4}$ and $A_{1}$. There is a small hybridization $0>V>t$ between lattice sites $A_{i}$ and $B_{i}$. The local Hubbard density-density interaction of strength $U\gg|t|$ is strongly repulsive. The connectivity and the coupling constants of the lattice models are chosen such that the lattice sites $B_{i}$ should be good realizations of spin-like orbitals.
@@
$H=$ $\sum_{i\sigma}\left(t\,a^{\dagger}_{i+1\sigma}a_{i\sigma}+V\,a^{\dagger}_{i\sigma}b_{i\sigma}+\text{h.c.}\right)$ (53)
@@
where $a^{\dagger}_{i\sigma}$ creates a fermion of spin $\sigma$ on lattice site $A_{i}$ and $b^{\dagger}_{i\sigma}$ creates a fermion of spin $\sigma$ on lattice site $B_{i}$. The lattice sites $A_{i}$ and $B_{i}$ form the initial basis of the system. We impose periodic boundary conditions for the lattice sites $A_{i}$. A graphic representation of the lattice model is displayed in Fig. 1. The coupling constants satisfy $U\gg|t|>|V|$ and $0>V>t$. The repulsive Hubbard interaction on lattice sites $B_{i}$ in combination with the weak hybridization between $B_{i}$ and $A_{i}$ places a significant energy penalty on $n_{B_{i}}\neq 1$. Low-energy eigenstates of the Hamiltonian should therefore satisfy $n_{B_{i}}\equiv 1$ and the lattice sites $B_{i}$ be considered good realizations of spin-like orbitals.
We have calculated the low energy spectrum of the Hamiltonian of a chain of $M=4$ unit cells containing $N=8$ fermions using a numerical diagonalization method. The reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ have been computed for the corresponding ground state $|\psi_{0}\rangle$. By definition [34], the eigenstates of $\rho^{(1)}$ form the natural (spin) orbital basis. We find eigenvalues $\langle n_{i}\rangle=1$ of $\rho^{(1)}$ for $N_{\phi_{i}}=2$ natural basis orbitals $\phi_{i}$. These are considered two natural spin orbitals.
Figure 2: Ground state expectation values of the local parity $\langle P_{q}\rangle_{0}$ of the initial basis orbitals $A_{i}$ and $B_{i}$, as well as (a) the natural (spin) orbitals basis and (b) the parity optimized basis for the spin-bath chain model described by eq. (53). (a) The dark grey line shows the local parities of the original basis orbitals $A_{i}$ and $B_{i}$ and the dashed orange line indicates the local parities of the natural basis orbitals. The blue circles display the expectation values for the electron density $\langle n_{q}\rangle_{0}$ in the natural orbital basis. In the natural orbital basis we find $N_{\phi_{q}}=2$ orbitals $\phi_{q}$ for which $\langle n_{q}\rangle_{0}=1$, but the respective values of the local parities $\langle P_{q}\rangle_{0}>0$ highlight that $|\delta n_{q}|\gg 0$. (b) The dashed orange line now displays the ground state local parities of the parity optimized basis orbitals. The blue circles show the expectation value for electron density $\langle n_{q}\rangle_{0}$ in these optimized basis orbitals. In the optimized basis we find $N_{\phi_{q}}=4$ orbitals $\phi_{q}$ with local parity $\langle P_{q}\rangle_{0}\simeq-1$. We also observe $\langle n_{q}\rangle_{0}=1$ for each orbital. The orbitals $\phi_{q\leq 3}$ of the optimized basis are considered spin-like. They coincide with the lattice sites $B_{i}$, but the ordering has been shuffled in the optimization procedure.